Yellow light from a sodium vapor lamp has a wavelength of
589 nm. What is the frequency of this electromagnetic wave?
1
Expert's answer
2020-10-25T18:28:05-0400

"f=\frac{c}{\lambda}=\frac{3\cdot10^8}{589\cdot10^{-9}}=509\cdot10^{12}(Hz)"